A 15.0 kg child slides down a 4.25 m playground slide that makes a 45o with the horizontal axis. While sliding, a 2.5 N force of friction acts on the child. What will the child’s speed be when she reaches the bottom of the slide?


1
Expert's answer
2021-02-19T10:31:17-0500

Let's apply the Newton's Second Law of Motion and find the acceleration of the child:


"mgsin\theta-F_{fr}=ma,""a=\dfrac{mgsin\theta-F_{fr}}{m},""a=\dfrac{15\ kg\cdot9.8\ \dfrac{m}{s^2}\cdot sin45^{\circ}-2.5\ N}{15\ kg}=6.76\ \dfrac{m}{s^2}."

Finally, we can find the child’s speed when she reaches the bottom of the slide from the kinematic equation:


"v^2=v_0^2+2ad,""v=\sqrt{v_0^2+2ad},""v=\sqrt{0+2\cdot6.76\ \dfrac{m}{s^2}\cdot4.25\ m}=7.58\ \dfrac{m}{s}."